OpenStudy (oraclethinktank):

x = y^2 - 1 2x - y = 4 Which of the following equations could be the result of using substitution to solve the system? 2y^2 - y - 5 = 0 2y^2 - y - 6 = 0 2y^2 - 2y - 5 = 0

OpenStudy (photon336):

\[x = y^{2}-1 \] see this right?

OpenStudy (oraclethinktank):

Yes

OpenStudy (photon336):

think of this as saying x = y^2-1 so wherever you see x you plug in y^2-1 where can you do this?

OpenStudy (oraclethinktank):

For the second equation in the system it would be 2(y^2-1) = y = 4

OpenStudy (photon336):

almost

OpenStudy (photon336):

\[2(y^{2}-1)-y = 4\]

OpenStudy (oraclethinktank):

Ah, I see the mistype.

OpenStudy (photon336):

know how to continue?

OpenStudy (oraclethinktank):

Not fully.

OpenStudy (photon336):

see that 2 out side ?

OpenStudy (photon336):

\[2(~~~~) \]

OpenStudy (oraclethinktank):

Yes.

OpenStudy (photon336):

\[2y^{2}-2-y = 4 \]

OpenStudy (photon336):

follow?

OpenStudy (photon336):

@OracleThinkTank

OpenStudy (photon336):

and also how would we continue based off of this

OpenStudy (oraclethinktank):

Okay.

OpenStudy (photon336):

yeah so how do you think we would continue?

OpenStudy (oraclethinktank):

I'm not too sure. I'm not the best with Algebra.

OpenStudy (photon336):

so the next step is we move all the terms to the same side.

OpenStudy (photon336):

so we have a 4 on one side right? and also, this is a quadratic equation

OpenStudy (oraclethinktank):

Okay.

OpenStudy (photon336):

\[2y^{2}-y-2 = 4 \]

OpenStudy (photon336):

now we subtract 4 from both sides. \[2y^{2}-y-2-4 = 0\] how would we simplify this?

OpenStudy (oraclethinktank):

3y^2 - 6 = 0?

OpenStudy (photon336):

\[2y^{2}-y-6 = 0 \] remember you can only combine like terms.

OpenStudy (oraclethinktank):

Oh, Thanks.